The following data show the yearly salaries (in £000s) of student union officers at eight universities.

University Salary

A 43    B 34 C 38 D 47 E 52 F 59 G 53 H 45

Question: What is the the standard deviation? (use 2 decimal places)
Answer #1

Answer)

First step is to find the mean

Mean = (sum of the observations)/(number of observations)

Mean = (43 + 34...)/8 = 46.375

Now we need to subtract mean from each and every observations and then we need to take the square and add the results

= (43 - 46.3755)^2 + (34 - 46.375)^2...

= 471.875

Now we need to divide 471.875 by n-1 that is by 7 and take the square root

S.d = √{471.875}/7} = 8.2104